OpenStudy (anonymous):

Which of the following represent the zeros of f(x) = 6x^3 – 31x^2 + 4x + 5 a. -5,1/3,1/2 b. 5, 1/3, 1/2 c. 5, 1/3, -1/2 d. 5, -1/3, 1/2

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

(x-5)(3x+1)(2x-1)=0 x=5 x=-1/3 x=1/2 5, –one–third, one–half